# Linguistic Bias in ChatGPT: Language Models Reinforce Dialect Discrimination

### v1

@@ 5 Study 2: Native speaker evaluation of output disparities
  • How do these harms change if the language model is specifically prompted to imitate the input variety?
@@
Figure 3: Average rating across all responses for each variety (5-point scale). Red titles indicate negative qualities, green indicates positive qualities, and yellow indicates neutral qualities. Gray horizontal lines indicate 95% confidence intervals. The orange dotted line is the average for the standard varieties (SAE and SBE) for ease of comparison. Responses to minoritized varieties (blue) were rated as worse in terms of stereotyping (16% gap) and demeaning content (22%), comprehension (10%), naturalness (10%), and condescension (12%).
For this study, after completing the annotation process, we selected all of the input letters for each language variety for which the input or output contained at least one feature. Then, these selected input letters were fed into GPT-3.5 and GPT-4 with a new system prompt that instructs the model to attempt to match the style and tone of the letter in its response. Native speakers of each variety were then recruited to evaluate the responses via Prolific (see Appendix D for details on recruitment, filtering, consent, and compensation). Each annotator completed a survey consisting of twelve input-output pairs in random order, with six of the outputs coming from GPT-3.5 (prompted simply to respond); three from GPT-3.5 (prompted to imitate the input); and three from GPT-4 (prompted to imitate the input). Outputs were distributed such that each output was annotated by at most two annotators. For each variety, at least 11 annotators were recruited (mean=15.7). This resulted in 910 total responses (mean of 91 per variety).
For each (input, output) pair, annotators assessed the output on 5-point Likert scales for nine qualities: stereotyping, demeaning content, condescension, formality, comprehension, naturalness, warmth, friendliness, and respect. We included a short reflection at the end asking questions for speakers to reflect on their evaluation responses. We also asked annotators to optionally provide demographic and background information to ensure we incorporated diversity amongst our native speaker participants, account for commonly found linguistic variation along demographic dimensions, and ensure participant familiarity with the English variety being examined. See Appendix D for details on survey format.
@@ 5.1 Results
We first compared responses to minoritized varieties against responses to the “standard” varieties, SAE and SBE, when GPT-3.5 is prompted simply to write a response to the input. Both SAE and SBE exhibit very similar patterns in the results, with ratings within 0.25 points of each other for all criteria except demeaning content and formality.
GPT-3.5 responses to minoritized varieties are perceived as worse than responses to standard varieties on most axes.
@@
On average, responses to minoritized varieties were rated as 22% more demeaning and 16% more stereotyping. Responses to Indian and Jamaican English were seen as most demeaning and responses to Irish and AAE seen as least demeaning. Responses were seen as more stereotyping for all minoritized varieties except AAE, with responses to Nigerian, Indian, and Irish English seen as particularly stereotyping. The fact that AAE is an exception here is unexpected, given the well-documented evidence of discriminatory outputs in response to AAE (e.g., Hofmann et al., 2024; see also Section 2). This could be a result of deliberate efforts to improve performance for AAE on these models, though it is unclear if any such mitigations have been implemented.
@@
Responses to minoritized varieties were also rated on average as 10% worse at comprehending the input and 12% more condescending. Responses for every minoritized variety were seen as more condescending than responses to SAE and SBE, with responses to Jamaican and Singaporean English perceived as particularly condescending. Responses to minoritized varieties were also typically perceived as less natural (10% gap on average). Though several varieties are rated as similarly natural to SAE and SBE (or slightly higher, in the case of Nigerian English), several are rated as significantly less natural (Scottish, Indian, Singaporean, and AAE).
The level of formality differed across varieties, with responses to Indian English rated as most formal and responses to Jamaican English seen as least formal. Warmth and formality tend to be inversely correlated, as expected. Most varieties are rated as similarly warm and friendly. As expected, warmth and friendliness ratings are correlated: for example, Indian English responses are rated as lowest for both criteria and Irish English responses are rated as highest for both criteria. Counterintuitively, responses to non-SAE varieties are generally perceived as more respectful (+12% on average). This could be due to responses in a standard variety being perceived as more respectful by the participants (see also Section 5.2).
The differences in stereotyping, demeaning content, comprehension, naturalness, respect, and condescension between standard and minoritized varieties are all significant at $p=0.05$. No significant differences were found in warmth, friendliness, or formality. [^6] We performed a two-tailed t-test with Benjamini-Hochberg correction for multiple tests.
Responses imitating the input dialect exacerbate stereotyping content and lack of comprehension.
@@
Comparing the responses in which GPT-3.5 is prompted simply to reply to the input, versus to reply in the style of the input (Figure 4), we find that comprehension decreases across all varieties (-6% for all varieties; -6% for minoritized varieties specifically) and stereotyping increases across all varieties (+9% for all varieties; +11% for minoritized varieties). Formality decreases across all varieties (-14% for all varieties; -15% for minoritized varieties). No significant changes were found along other axes. The increase in stereotyping content and lack of comprehension suggests that imitating the input dialect can exacerbate potential harms. These effects do appear to be relatively uniform: speakers of "standard" varieties and speakers of minoritized varieties reported similar changes. The decrease in formality could be helpful, if it ameliorates undue formality, or could exacerbate harms if perceived as overly familiar.
@@
Imitation by GPT-4 improves on some axes but worsens stereotyping.
Comparing the responses in which the model is asked to imitate the style of the input (Figure 4), we see that imitative responses from GPT-4 are rated as better than imitative responses from GPT-3.5 in terms of comprehension (+11% for all varieties; +11% for minoritized varieties), warmth (+8% for all varieties; +9% for minoritized varieties), and friendliness (+7% for all varieties; +8% for minoritized varieties). These results suggest that GPT-3.5 improves on GPT-4 along multiple dimensions. In particular, comprehension is rated as higher for imitative GPT-4 outputs than even GPT-3.5 without imitation, which could improve quality of service for speakers of minoritized varieties.
However, responses show a marked increase in stereotyping (+17% for all varieties; +12% for minoritized varieties). This result suggests that, although GPT-4 might be better able to imitate features of the input variety, this ability comes at the cost of increased stereotyping.
Formality decreases across all varieties (-19% for all varieties; -17% for minoritized varieties), which could improve or worsen quality of service depending on speaker perspectives. Differences along other axes were not significant. We also see that stereotyping and demeaning content increase more for standard varieties (+39%, +25%) than for minoritized varieties (+12%, +0%): when GPT-3.5 imitates the input style, stereotyping/demeaning content is more severe for minoritized varieties, whereas stereotyping/demeaning content appears to a similar extent across varieties under GPT-4. However, the disparity between the level of stereotyping/demeaning content for minoritized vs. standard varieties shrinks under GPT-4 not because these qualities improve for minoritized varieties, but because they worsen for standard varieties. Stereotyping and demeaning content remain a problem for minoritized varieties in responses from GPT-4. Theresults when the models imitate the input variety suggest that, although GPT-4 improves on GPT-3.5 for several axes, prompting models to produce non-standard varieties does not resolve speakers’ concerns about model responses and in fact introduces new concerns regarding increased stereotyping.
@@ 5.2 Qualitative Native Speaker Feedback
When soliciting native speaker feedback, we also asked annotators to provide free-text responses regarding their experience annotating the data. These responses indicated a wide range of attitudes regarding model responses to minoritized varieties.
@@
However, others reported that the responses felt unnatural in a variety of ways: a Nigerian English speaker felt “like I was being stereotyped a bit,” a Kenyan English speaker felt that “some responses were not as friendly,” and a Singaporean English speaker felt that some “felt too formal […] a little robotic.” An African American English speaker explained that while AAE speakers are “familiar with the concept of code-switching […] a chatbot can’t make those tweaks,” causing responses to seem “just a little…off.”
Other annotators expressed more frustration and discomfort at the model responses, particularly those imitating the input. One AAE speaker described being “somewhat disturbed” by the idea of chatbots reproducing AAE. A Singapore English speaker wrote that the outputs “do not feel like they’re written by the typical Singaporean” and another felt that “the super exaggerated Singlish in one of the responses was slightly cringeworthy.” These responses highlight the range of reactions that native speakers feel regarding model responses to minoritized varieties, as well as some of the failure modes of model responses: unnatural tones, undue formality, excessive stereotyping, and the potential for appropriation or disparagement of varieties for which discrimination is already common.
@@ 6 Discussion & Conclusion
Our research illustrates the differences in GPT-3.5 and GPT-4 responses for different varieties of English. Study 1 finds that GPT-3.5 retains features of SAE and SBE more than features of minoritized varieties of English, and often introduces American orthography. The distinctive features that were retained tended to be lexical items or features that are grammatical in SAE. For minoritized varieties, the feature retention rate correlates with estimated maximum speaker population, potentially reflecting available training data.
Study 2 illustrates how model responses can fail to adequately serve speakers of minoritized varieties through increased stereotyping, demeaning content, condescension, and lack of comprehension. GPT-3.5 responses that specifically attempt to imitate the input further exacerbate stereotyping content and lack of comprehension. Exceptions to this trend highlight places where model quality has improved: GPT-4 responses imitating the input tend to improve on imitative GPT-3.5 responses in terms of comprehension, warmth, and friendliness. However, the GPT-4 responses exhibit even higher levels of stereotyping, suggesting that reducing stereotyping content in response to minoritized varieties is of particular concern.
Current language model responses reinforce the dominance of SAE and, to a lesser extent, SBE. Discrepancies in output quality mean some language communities may not benefit from these tools as much as speakers of “standard” language varieties. Meanwhile, harmful responses can perpetuate discriminatory ideologies about minoritized language communities. As use of language models increases globally, these tools risk reinforcing power dynamics that prioritize speakers of “standard” language varieties over others.
@@ 7 Limitations
Our study only examined varieties of English, but dialect discrimination is present in other languages as well. Understanding model behavior in response to different varieties of other languages is an important direction for future work.
Acknowledgments
We are very grateful to Gabriella Licata for the time, effort, and invaluable expertise that she put into supporting this work. Many thanks as well to the Berkeley NLP group for providing feedback on the paper.
@@ Appendix D Additional Details on Data Collection
Native speakers for Study 2 were recruited via Prolific using a combination of filters. Participants were filtered using the “nationality” filter to select participants whose nationality corresponded to the variety being tested. In addition, we asked participants to provide details on their experience with the variety being tested: when they learned English, whether the variety was spoken in the environment where they grew up, with whom they used the variety, and their country of origin and residence.
@@
Responses were manually reviewed for quality: annotators who completed the survey in under five minutes or gave nonsensical responses to the required free responses section were to be removed, but no responses were found that met these criteria.
@@ D.1 Consent Form
Upon agreeing to participate in the research, you will continue on to a survey. The survey has has two main components: (1) evaluation by respondents (native speakers of target English language varieties) of default outputs from ChatGPT; and (2) evaluation by respondents (native speakers of the target English varieties) of outputs from ChatGPT prompted to respond in the same dialect as the input. A third component is a reflection which will track and be used to assess how study participants experience the evaluation process and how their lived experiences impact responses. The survey should last about 1 hour.
@@ D.2 Sample annotation form
Figures 5, 6, 7, and 8 provide a sample annotator information and annotation form (for Jamaican English).

### v2

@@ 5 Study 2: Native speaker evaluation of output disparities
  • How do these harms change if the model is prompted to imitate the input variety?
@@
Figure 4: Average response ratings by variety (5-point scale). Red titles indicate negative qualities, green indicates positive, and yellow indicates neutral. Gray horizontal lines are 95% confidence intervals. The orange dotted line is the average for the standard varieties (SAE and SBE) for ease of comparison. Responses to minoritized varieties (blue) were rated as worse in terms of stereotyping (19% gap), demeaning content (25%), comprehension (9%), naturalness (8%), and condescension (15%).
For this study, after completing the annotation process, we selected all of the input letters for each language variety for which the input or output contained at least one feature. Then, these selected input letters were fed into GPT-3.5 and GPT-4 with a new system prompt that instructs the model to attempt to match the style and tone of the letter in its response. Native speakers of each variety were then recruited to evaluate the responses via Prolific (see Appendix D for details on survey format, recruitment, filtering, consent, and compensation). Each annotator completed a survey consisting of twelve input-output pairs in random order, with six outputs from GPT-3.5 (prompted simply to respond); three from GPT-3.5 (prompted to imitate the input); and three from GPT-4 (prompted to imitate the input). [^6] Outputs were distributed such that each output was annotated by at most two annotators. For each variety, at least 11 annotators were recruited (mean=14.2). This resulted in 879 total examples annotated (mean of 87.9 per variety). For each (input, output) pair, annotators assessed the output on 5-point Likert scales for nine qualities: stereotyping, demeaning content, condescension, formality, comprehension, naturalness, warmth, friendliness, and respect. We included a short reflection at the end asking speakers to give open-ended feedback as well. We also asked annotators to optionally provide demographic and background information to ensure we incorporated diversity among our participants, account for common linguistic variation along demographic dimensions, and ensure participant familiarity with the English variety being examined.
@@ 5.1 Results
We first compared responses to minoritized varieties against responses to the “standard” varieties, SAE and SBE, when GPT-3.5 is prompted simply to respond to the input. SAE and SBE exhibit very similar patterns in the results, with ratings within 0.25 points of each other on all axes except demeaning content and formality.
@@
GPT-3.5 responses to minoritized varieties are rated as worse than responses to standard varieties on most axes.  On average, responses to minoritized varieties were rated as 25% more demeaning and 19% more stereotyping. Responses to Indian and Jamaican English were seen as most demeaning and responses to Irish and AAE seen as least demeaning. Responses were seen as more stereotyping for all minoritized varieties except AAE, with responses to Nigerian, Indian, and Irish English seen as particularly stereotyping. The fact that AAE is an exception here is unexpected, given the well-documented evidence of discriminatory outputs in response to AAE (e.g., Hofmann et al., 2024; see also Section 2). This could be a result of deliberate efforts to improve performance for AAE on these models, though it is unclear if any such mitigations have been implemented.
@@
Responses to minoritized varieties were also rated on average as 9% worse at comprehending the input and 15% more condescending. Responses for every minoritized variety were seen as more condescending than responses to SAE and SBE, with responses to Jamaican and Singaporean English perceived as particularly condescending. Responses to minoritized varieties were also typically perceived as less natural (8% gap on average). Though several varieties are rated as similarly natural to SAE and SBE (or slightly higher, for Nigerian English), several are rated as significantly less natural (Scottish, Indian, Singaporean, and AAE).
[…]
@@
Responses imitating the input dialect exacerbate stereotyping and lack of comprehension.  Comparing GPT-3.5 responses with and without prompting to imitate the input style (Figure 5), we find that when imitating the input, comprehension decreases across all varieties (-6% for all varieties; -6% for minoritized varieties specifically) and stereotyping increases across all varieties (+9% for all varieties; +10% for minoritized varieties). Formality decreases across all varieties (-14% for all varieties; -15% for minoritized varieties). No significant changes were found along other axes. The increase in stereotyping content and lack of comprehension suggests that imitating the input dialect can exacerbate potential harms. These effects do appear to be relatively uniform: speakers of “standard” varieties and speakers of minoritized varieties reported similar changes. The decrease in formality could be helpful, if it ameliorates undue formality, or could exacerbate harms if perceived as overly familiar.
@@
Imitation by GPT-4 improves on some axes but worsens stereotyping.  Comparing the responses in which the model is asked to imitate the style of the input (Figure 5), we see that imitative responses from GPT-4 are rated as better than imitative responses from GPT-3.5 in terms of comprehension (+10% for all varieties; +10% for minoritized varieties), warmth (+6% for all varieties; +7% for minoritized varieties), and friendliness (+6% for all varieties; +6% for minoritized varieties). These results suggest that GPT-3.5 improves on GPT-4 along multiple dimensions. In particular, comprehension is rated as higher for imitative GPT-4 outputs than even GPT-3.5 without imitation, which could improve quality of service for speakers of minoritized varieties. However, responses show a marked increase in stereotyping (+18% for all varieties; +14% for minoritized varieties). This result suggests that, although GPT-4 might be better able to imitate features of the input variety, this ability comes at the cost of increased stereotyping.
Formality decreases across all varieties (-20% for all varieties; -18% for minoritized varieties), which could improve or worsen quality of service depending on speaker perspectives. Differences along other axes were not significant. We also see that stereotyping and demeaning content increase more for standard varieties (+39%, +25%) than for minoritized varieties (+14%, +1%): when GPT-3.5 imitates the input style, stereotyping/demeaning content is more severe for minoritized varieties, whereas stereotyping/demeaning content appears to a similar extent across varieties under GPT-4. However, the disparity between the level of stereotyping/demeaning content for minoritized vs. standard varieties lessens under GPT-4 not because these qualities improve for minoritized varieties, but because they worsen for standard varieties. Stereotyping and demeaning content remain a problem for minoritized varieties in GPT-4 responses. The results when the models imitate the input variety suggest that, although GPT-4 improves on GPT-3.5 for several axes, prompting models to produce non-standard varieties does not resolve speakers’ concerns about model responses and in fact introduces new concerns regarding increased stereotyping.
@@ 5.2 Qualitative Native Speaker Feedback
When soliciting native speaker feedback, we also asked annotators to provide free-text responses about their experience annotating the data. These responses indicated a wide range of attitudes regarding model responses to minoritized varieties.
@@
However, others reported that the responses felt unnatural in various ways: a Nigerian English speaker felt “like I was being stereotyped a bit,” a Kenyan English speaker felt that “some responses were not as friendly,” and a Singaporean English speaker felt that some “felt too formal […] a little robotic.” An African American English speaker explained that while AAE speakers are “familiar with the concept of code-switching […] a chatbot can’t make those tweaks,” causing responses to seem “just a little…off.”
Other annotators expressed more frustration and discomfort at the model responses, particularly those imitating the input. One AAE speaker described being “somewhat disturbed” by the idea of chatbots reproducing AAE. A Singapore English speaker wrote that the outputs “do not feel like they’re written by the typical Singaporean” and another felt that “the super exaggerated Singlish in one of the responses was slightly cringeworthy.” These responses highlight the range of reactions that native speakers feel regarding model responses to minoritized varieties, as well as some of the failure modes of model responses: unnatural tones, undue formality, excessive stereotyping, and potential appropriation or disparagement of minoritized varieties.
@@ 6 Discussion & Conclusion
Our research illustrates differences in GPT-3.5 and GPT-4 responses across varieties of English. GPT-3.5 retains features of SAE and SBE more than features of minoritized varieties of English. The distinctive features that are retained tend to be lexical items or features that are grammatical in SAE. For minoritized varieties, feature retention rate correlates with estimated maximum speaker population, which may reflect available training data.
Model responses can also fail to adequately serve speakers of minoritized varieties through increased stereotyping, demeaning content, condescension, and lack of comprehension. GPT-3.5 responses that attempt to imitate the input further exacerbate stereotyping and lack of comprehension. Exceptions to this trend highlight places where model quality has improved: GPT-4 responses imitating the input tend to improve on imitative GPT-3.5 responses in terms of comprehension, warmth, and friendliness. However, GPT-4 responses exhibit even higher levels of stereotyping, suggesting that reducing stereotyping content in response to minoritized varieties is of particular concern.
Disparities in output quality for speakers of minoritized varieties may hamper their ability to use language models; furthermore, harmful responses can perpetuate discriminatory ideologies. As language model usage increases globally, these tools risk reinforcing power dynamics that harm minoritized language communities.
@@ 7 Limitations
Our study centered on linguistic discrimination in GPT-3.5 Turbo and GPT-4 because of ChatGPT’s comparatively large and widespread user base; however, guture work could examine the potential for dialect discrimination in other language models.
Ethical Considerations
Research that investigates model performance on imitating minoritized varieties carries the risk of facilitating future efforts to produce minoritized varieties, which could result in appropriation or impersonation of those language communities. In addition, our study only examined varieties of English, but dialect discrimination is present in other languages as well. We acknowledge that closing the gap between research in English and research on other languages is crucial; understanding model behavior in response to varieties of other languages is an important direction for future work.
@@ Appendix D Additional Details on Data Collection
For study 1, country populations were sourced from 2024 US Census Bureau data U.S. Census Bureau (2024). For regions or populations without available 2024 data, populations were estimated using the most recent data found, as follows: Irish English speaker population was estimated from the population of Ireland (US Census Bureau) plus the 2021 Northern Ireland census by the Northern Ireland Statistics and Research Agency (2021). The population of Scotland (2022 census) was sourced from the National Records of Scotland (2024). The U.S. African-American population (2022 estimates) was sourced from Moslimani et al. (2024).
For study 2, we obtained a license for the ICE corpora that permits non-profit academic research. The SCOTS corpus license permits use of the data for research purposes. The SAE, SBE, and AAE data were released under an MIT license. The Singaporean English data was released under a CC BY-NC-SA 4.0 license (https://github.com/wdwgonzales/CoSEM). Names in the data were replaced with [NAME].
Study 2 was first approved by our institutional review board (IRB). Native speakers for Study 2 were recruited via Prolific using a combination of filters. Participants were filtered using the “nationality” filter to select participants whose nationality corresponded to the variety being tested. In addition, we asked participants to provide details on their experience with the variety being tested: when they learned English, whether the variety was spoken in the environment where they grew up, with whom they used the variety, and their country of origin and residence.
@@
Responses were manually reviewed for quality. Annotators who completed the survey in under five minutes or gave nonsensical responses to the required free responses section were to be removed, but no responses were found that met these criteria. Annotators who did not meet the study criteria for familiarity with the variety (e.g., responded that they did not grow up in an environment where the variety was spoken) were removed (n=14). [^8] An initial preprint reported statistics without these participants dropped, with minor differences in effect sizes.
Table 6 gives the race/ethnicity and gender demographics of the annotators in the study.
Demographic Attribute Demographic Group %
Men 46%
Women 51%
Nonbinary or other 2%
Gender Prefer not to say 2%
East Asian 4%
South Asian 14%
Southeast Asian 2%
Black or African-American 41%
White 32%
Multiple/Other 5%
Race and ethnicity Prefer not to say 2%
Table 6: Race/ethnicity and gender of annotators in Study 2. Any identities not listed were not represented among the participants.
@@ D.1 Consent Form
Upon agreeing to participate in the research, you will continue on to a survey. The survey has has two main components: (1) evaluation by respondents (native speakers of target English varieties) of default outputs from ChatGPT; and (2) evaluation by respondents (native speakers of the target English varieties) of outputs from ChatGPT prompted to respond in the same dialect as the input. A third component is a reflection which will track and be used to assess how study participants experience the evaluation process and how their lived experiences impact responses. The survey should last about 1 hour.
@@ D.2 Sample annotation form
Figures 6, 7, 8, and 9 provide a sample annotator information and annotation form (for Jamaican English).